# Can It Edit? Evaluating the Ability of Large Language Models to Follow Code Editing Instructions

### 05 Mar 2024 v4

@@ 4. Fine-tuning
For our ablation studies, we focus on the variant with 6.7 billion parameters. This model offers an ideal balance between size and performance, allowing us to extrapolate results to larger models with more parameters without the need for extensive training, which would be resource-intensive and time-consuming. Following the most efficient training strategy identified, we also fine-tune the 1.3b and 33b models to evaluate the impact of model size on code editing performance. Our fine-tuned models are referred to as EditCoder.
4.1. Training Data
Figure 3. Sunburst plot of the top 20 most frequent initial verbs, with their corresponding top 10 root nouns, in the commit messages of Commits2023FT.
(a) EditPackFT
(b) Commits2023FT
Figure 4. The distribution of the number of lines in the ‘before’ and ‘after’ code segments in the EditPackFT and Commits2023FT datasets. The 99th percentile is removed for clarity.
@@ 4.1. Training Data
Code Segments Mean Std. Dev. Mean Std. Dev.
Lines of Code 29.2 13.7 119.3 75.9
Levenshtein Distance 197.1 260.6 406.6 631.2
Commit Message Mean Std. Dev. Mean Std. Dev.
Tokens 10.1 4.6 23.1 35.2
Table 2. Training dataset statistics for EditPackFT and Commits2023FT
@@
We created the EditPackFT dataset by further filtering the Python split of the CommitPackFT dataset (Muennighoff et al., 2023), which was used to train OctoCoder. CommitPack is an extensive dataset comprising 4TB of permissively licensed commits from a 2016 GitHub snapshot across various programming languages. CommitPackFT is a subset of CommitPack, filtered for it to be amenable to instruction-tune Code LLMs. The primary criterion for CommitPackFT’s selection involved retaining commits whose messages begin with an imperative verb, mirroring the typical structure of natural language instructions. We apply a series of additional filtering steps, which make the dataset more suitable for code editing. We remove any item that passes any of the following predicates:
  1. (1)
@@
Originally, the dataset contained 56,025 commits, and after applying the filtering steps, we are left with 22,602 commits. As shown by Figure 3(a) and Table 2, the mean number of lines in the ‘before’ and ‘after’ code segments is $29.2$. We further analyzed the original CommitPackFT dataset, ensure that our filtering wasn’t the cause of the short code segments, and find that the mean number of lines is similar, with a mean of $28.8$. This distribution may be suitable for small-scale code editing tasks. We also analyze the distribution of the commit message lengths, and find that the mean token count is $10.1$, which is quite low.
@@
To address the limitations of EditPackFT, we developed the Commits2023FT dataset. This dataset consists of 416,792 Python file changes from commits in permissively licensed GitHub repositories, and is named Commits2023. Our objective is to create a dataset akin to CommitPackFT, but with more recent data and a more diverse example length distribution. We employed the same filters on this dataset as used for EditPackFT, and also applied the initial filters from CommitPackFT, which excludes any commit lacking a message that starts with an imperative verb, or a message that is not within the 10 to 1000 character length range. Additionally, we only retain one file from multi-file commits to avoid exact duplicate commit messages in our dataset (Lee et al., 2022). After this filtering process, we obtain a dataset comprising 24,129 Python file changes. Figure 3(b) presents a broader distribution of the number of lines in the ‘before’ and ‘after’ code segments, with an average of $119.3$. Figure 3 illustrates a sunburst plot of the most frequent initial verbs in the commit messages of Commits2023FT, along with their corresponding root nouns. This set of verbs is slightly more varied than those in EditPackFT, featuring $199$ unique verbs in comparison to $184$. Furthermore, the token count distribution of the commit messages is two times higher and much more varied than that of EditPackFT, with a mean of $23.1$ and a standard deviation of $35.2$.
@@
For ablation analysis, we generated two additional datasets: Commits2023Raw25k and Commits2023FT+EditPackFT. Commits2023Raw25k consists of a random selection of 25,000 commits from Commits2023, we use this dataset to assess the impact of the filtering process on the final dataset. Commits2023FT+EditPackFT represents the combined dataset of Commits2023FT and EditPackFT. In our evaluation, we find that the combination of Commits2023FT and EditPackFT yields the best results by a significant margin (Section 5.2), and thus we train our final models on this dataset. We believe that these results are due to the increased volume of data, and the more varied length distributions.
4.2. Training Tools and Configuration
## Code Before:{before}## Instruction:{instruction}## Code After:{after}
Figure 5. Prompt format for EditCoder.
For training all of our EditCoder models, we utilize a fine-tuning pipeline based on the HuggingFace Transformers library (Wolf et al., 2020). Additionally, we utilize DeepSpeed ZeRO 3 (Rajbhandari et al., 2020) to efficiently shard the model and dataset across multiple GPUs. We also use FlashAttention 2 (Dao, 2023) to speed up training on large context window sizes. All of our models are trained on a single machine equipped with 8 NVIDIA A100 (80GB) GPUs. The effective micro-batch size is set at $32$ ($4$ gradient accumulation steps, with a single batch per GPU). We employ a learning rate of $2\times 10^{-5}$ with linear decay and $10$ warmup steps. All models underwent training for $8$ epochs, with a constant, unpadded context window of $8192$ tokens. Prior to training, we shuffled the dataset randomly and deduplicated [^1] Deduplication, achieved by concatenating the ‘before’ and ‘after’ code segments, helps mitigate overfitting to specific training examples (Lee et al., 2022). it following the method outlined by Li et al. (2023a). This process combines MinHash (Broder, 2000) and Locality Sensitive Hashing (LSH) (Leskovec et al., 2014). We format the training data as a prompt, with the ‘before’ code segment followed by the ‘instruction’ and the ‘after’ code segment, as show in Figure 5.
5. Evaluation
Model pass@1 ExcessCode
Name Size Descriptive Lazy Descriptive Lazy
@@ 5. Evaluation
GPT-4 61.85 54.72 0.39 0.0
GPT-3.5-Turbo 58.98 46.48 1.74 1.57
Unknown-data Open Models
Deepseek-Coder-Instruct 33b 53.06 43.89 1.53 1.26
Deepseek-Coder-Instruct 6.7b 33.89 33.61 0.19 0.44
Deepseek-Coder-Instruct 1.3b 25.83 18.33 1.02 0.44
@@
EditCoder 33b 50.19 40.21 0.48 0.0
EditCoder 6.7b 52.59 38.46 0.4 0.03
EditCoder 1.3b 30.5 25.06 2.04 1.89
CodeLlama-Instruct 34b 35.0 26.76 0.37 0.67
CodeLlama-Instruct 13b 28.33 20.19 2.52 0.0
CodeLlama-Instruct 7b 33.89 27.04 0.12 0.18
Deepseek-Coder-Base 33b 32.37 23.26 0.39 0.52
Deepseek-Coder-Base 6.7b 28.43 22.95 2.0 0.0
Deepseek-Coder-Base 1.3b 0.37 1.11 1.0 24.0
StarCoder 15b 37.31 29.16 0.67 1.23
StarCoderBase 15b 38.24 26.38 1.70 1.48
StarCoderBase 7b 40.64 25.83 0.92 0.15
StarCoderBase 3b 19.62 12.78 1.13 0.0
StarCoderBase 1b 8.70 9.07 2.0 0.0
OctoCoder 15b 31.46 25.69 0.23 0.17
Table 3. Evaluation results of close and open-access models on CanItEdit. We report the pass@1 and ExcessCode metrics for both the descriptive and lazy prompts as well as the size of the model if available.
In this section, we evaluate the performance of various open and closed-sourced models on the CanItEdit benchmark, as well our fine-tuned models.
@@
We run the open-access models using HuggingFace Transformers (Wolf et al., 2020) and vLLM (Kwon et al., 2023). We use the following hyperparameters for all inference experiments: batch size $100$, $8192$ maximum new tokens, temperature $0.2$, and top-$p$ sampling cutoff of $0.95$. We run all tests in a Docker container to mitigate the risk of malicious code execution.
@@
We evaluate several state of the art models with varying sizes, and also fine-tune some models to build EditCoder. We group the models into three categories: open, closed-sourced, and unknown-data open models, where the latter are open models fine-tuned on an unknown dataset, which makes it difficult to compare to other models. The full list of models and their sizes appears in Table 3.
Finally, we were careful in formatting each benchmark problem to use prompt formats that the models’ developers recommend. The specific formats appear in Appendix A.
5.1. Evaluation Metrics
We employed two metrics to assess the performance of different models: one for functional correctness and another for the conciseness of the code edits.
  • pass@1 calculates the average fraction of successful completions per problem in CanItEdit, where success is defined as a completion passing all unit tests. Following Cassano et al. (2023b), we generated 20 completions per problem.
  • Besides functional correctness, we assess the conciseness of model-generated code edits using the ExcessCode metric. This metric evaluates the presence of unexecuted code in successful completions by calculating the percentage of superfluous code, as indicated by the percentage line coverage in the generated code. We calculate this metric by averaging the median line coverage for passing completions across all problems, omitting those with no successful completions.
5.2. Results with Existing Models
Training Dataset pass@1 ExcessCode
Name #Tokens #Items Descriptive Lazy Descriptive Lazy
Commits2023FT+EditPackFT 74M 46,274 52.59 38.46 0.4 0.03
Commits2023FT 62M 24,129 48.31 35.96 0.47 0
Commits2023Raw25k 62M 25,000 47.41 36.65 0.47 0
EditPackFT 12M 22,602 48.15 36.96 0.75 0
Table 4. Ablation results of training Deepseek-Coder-6.7b-Base on different datasets and evaluating on CanItEdit. We show the total number of tokens and items in each dataset, as well as the pass@1 and ExcessCode metrics for both the descriptive and lazy prompts. The reported sizes of the datasets are after deduplication.
@@ 5.2. Results with Existing Models
Larger models are better at editing; small models generate more excess code. Generally, model size correlates positively with pass@1 and negatively with ExcessCode. This indicates that larger models are more adept at precise functionality addition.
Models pre-trained on commits are better at code editing. Of the open models, the StarCoder model family is unique because it is pre-trained on a sample of GitHub commits (Li et al., 2023a), and we use the StarCoder commit data format when we evaluate the StarCoder models. We find that StarCoder models are significantly better on our benchmark than the pre-trained DeepSeek and Code Llama models, despite the fact that the latter two models outperform StarCoder on code generation (Guo et al., 2024).
Models are generally better at following descriptive instructions than lazy instructions. Models generally perform better with descriptive instructions, likely because these provide more specific code details. However, some smaller models like Deepseek-Coder-Base-1.3b and StarCoderBase-1b perform better with lazy instructions, possibly due to their limited capacity to process longer detailed instructions. Detailed statistics are available in Table 1.
Closed and unknown-data models outperform open models. The comparison between CodeLlama-Instruct, a generic instruction-following code generation model, and GPT-4, a broad instruction-following model, highlights the performance gap between open and closed-sourced models  (Roziere et al., 2023; OpenAI, 2023a). In terms of pass@1, GPT-4 outperforms CodeLlama-Instruct-34b by $26.85\%$ and $14.51\%$ for descriptive and lazy instructions, respectively, confirming the significant gap in instructional code editing abilities between state-of-the-art open source and proprietary models.
5.3. Results after Fine-Tuning on Commits
@@
In addition to evaluating existing open models, we also fine-tuned pre-trained DeepSeek models (section 4) to build EditCoder, which we now evaluate.
@@
Optimal Dataset: Commits2023FT+EditPackFT. In finding the best training dataset for Deepseek-Coder-6.7b-Base, the base model for EditCoder, various ablation datasets were tested. Results in Table 4 show Commits2023FT+EditPackFT as the top performer for both descriptive and lazy instructions. The dataset’s larger size and diverse data types, including varied commits, edits, and instructions, likely contribute to its superior performance.
Fine-tuning on open commits can significantly improve code editing performance. EditCoder-6.7b surpasses all open models, showing an $11.95\%$ increase in pass@1 and a notable decrease in ExcessCode compared to StarCoderBase-7b for descriptive instructions. Despite being smaller than the EditCoder-33b, its performance advantage is attributed to the smaller training dataset size (46,274 items), which might lead to overfitting in larger models. EditCoder-33b, however, performs better with lazy instructions, suggesting its efficiency in inferring user intent from minimal instructions. We see a similar trend with StarCoderBase-7b performing better than its 15b variant on descriptive instructions. Being both trained on commit data, might suggest a connection between model size, training data nature, and code editing performance. This observation highlights how targeted fine-tuning on code editing datasets distinctively boosts performance, demonstrating the unique demands of instructional code editing over other programming tasks.
6. Conclusion

### 20 Mar 2024 v5

@@ 4 Fine-Tuning
For our ablation studies, we focus on the model with 6.7 billion parameters, which offers an ideal balance between size and performance. This allows us to extrapolate results to larger models with more parameters without the need for extensive training. Following the most performant training strategy identified, we also fine-tune the 1.3b and 33b models to evaluate the impact of model size on code editing performance. Our fine-tuned models are referred to as EditCoder. These models have been full-parameter fine-tuned by calculating the loss on only the ‘after’ code segment. Appendix B provides further details and experiments on the training process.
@@
Code Segments (Mean $\pm$ Std. Dev.)
Lines of Code $29.2\pm 13.7$ $119.3\pm 75.9$
Levenshtein Distance $197.1\pm 260.6$ $406.6\pm 631.2$
Commit Messages (Mean $\pm$ Std. Dev.)
Tokens $10.1\pm 4.6$ $23.1\pm 35.2$
Table 2: Training dataset statistics for EditPackFT and Commits2023FT
@@
(a) EditPackFT
(b) Commits2023FT
Figure 2: The distribution of the number of lines in the ‘before’ and ‘after’ code segments in the EditPackFT and Commits2023FT datasets. The 99th percentile is removed for clarity.
@@
We created the EditPackFT dataset by further filtering the Python split of the CommitPackFT dataset (Muennighoff et al., 2023). CommitPack is an extensive dataset comprising 4TB of permissively licensed commits from a 2016 GitHub snapshot across various programming languages. CommitPackFT is a subset of CommitPack, filtered to be amenable for instruction-tuning Code LLMs. The primary criterion for CommitPackFT’s selection involved retaining commits whose messages begin with an imperative verb, mirroring the typical structure of natural language instructions. We apply a series of additional filtering steps, which make the dataset more suitable for code editing. We remove any item that pass any of the following predicates:
  1. 1.
@@
Originally, the dataset contained 56,025 commits, and after applying the filtering steps, we are left with 22,602. As shown by Figure 1(a) and Table 2, the mean number of lines in the ‘before’ and ‘after’ code segments is $29.2$. The mean Levenshtein distance between the ‘before’ and ‘after’ code segments is $197.1$ characters, indicating that the changes are relatively small. We also analyze the distribution of the commit message lengths, and find that the mean token count is $10.1$, which is quite low. We further analyzed the original CommitPackFT dataset to ensure that these findings weren’t artifacts of our filtering strategy: the full dataset has similar statistics.
@@
Figure 3: Sunburst plot of the top 20 most frequent initial verbs, with their corresponding top 10 root nouns, in the commit messages of Commits2023FT.
To address the limitations of EditPackFT, we developed the Commits2023FT dataset. This dataset is filtered from 416,792 Python file changes gathered from commits in permissively licensed GitHub repositories. We name the unfiltered dataset Commits2023. Our objective is to create a dataset akin to CommitPackFT, but with more recent data and a more diverse example length distribution. We employed the same filters on this dataset as used for EditPackFT, and also applied the initial filters from CommitPackFT, which includes only commits with messages that start with an imperative verb followed by at least a noun. Additionally, we only retain one file from multi-file commits to avoid exact duplicate commit messages in our dataset. After this filtering process, we obtain a dataset comprising 24,129 Python file changes. Figure 1(b) presents a broader distribution of the number of lines in the ‘before’ and ‘after’ code segments, with an average of $119.3$. We also observe a much larger change distribution, with a mean Levenshtein distance of $406.6$, signaling that this is not only a dataset with larger code segments, but also contains more varied changes. Figure 3 illustrates a sunburst plot of the most frequent initial verbs in the commit messages of Commits2023FT, along with their corresponding root nouns. This set of verbs is slightly more varied than those in EditPackFT, featuring $199$ unique verbs in comparison to $184$. Furthermore, the token count distribution of the commit messages is twice as high and much more varied than that of EditPackFT, with a mean of $23.1$ and a standard deviation of $35.2$.
@@
For ablation analysis, we generated two additional datasets: Commits2023Raw25k and Commits2023FT+EditPackFT. Commits2023Raw25k consists of a random selection of 25,000 commits from Commits2023. We use this dataset to assess the impact of the filtering process on the final dataset. Commits2023FT+EditPackFT represents the combined dataset of Commits2023FT and EditPackFT. We find that the combination of Commits2023FT and EditPackFT yields the best results by a significant margin (§ 5.2), and thus we train our final models on this dataset. We believe that these results are due to the increased amount of data and the expanded length distributions.
5 Evaluation
Model Descriptive Lazy
Name Size pass@1 ExcessCode pass@1 ExcessCode
@@ 5 Evaluation
GPT-4 63.33 0.15 $\pm$ 0.09 51.95 0.14 $\pm$ 0.10
GPT-3.5-Turbo 48.14 0.47 $\pm$ 0.34 42.71 0.00 $\pm$ 0.00
@@
CodeLlama-Instruct 70b 45.05 0.28 $\pm$ 0.15 37.52 0.02 $\pm$ 0.02
Mixtral-Instruct 8x7b 30.10 0.40 $\pm$ 0.16 24.90 0.01 $\pm$ 0.01
EditCoder 33b 55.90 0.33 $\pm$ 0.21 42.33 0.27 $\pm$ 0.24
DeepSeekCoder-Instruct 33b 49.78 0.36 $\pm$ 0.24 38.94 0.51 $\pm$ 0.34
DeepSeekCoder-Base 33b 47.71 0.53 $\pm$ 0.24 34.71 0.62 $\pm$ 0.41
CodeLlama-Instruct 34b 30.63 0.33 $\pm$ 0.21 24.15 0.18 $\pm$ 0.14
StarCoder2 15b 41.95 0.36 $\pm$ 0.20 31.48 0.04 $\pm$ 0.04
StarCoder 15b 37.10 0.56 $\pm$ 0.28 27.62 0.42 $\pm$ 0.34
OctoCoder 15b 34.43 0.12 $\pm$ 0.07 25.95 0.07 $\pm$ 0.07
CodeLlama-Instruct 13b 26.90 0.90 $\pm$ 0.68 16.89 0.42 $\pm$ 0.41
EditCoder 6.7b 48.33 0.36 $\pm$ 0.17 39.29 0.32 $\pm$ 0.25
DeepSeekCoder-Instruct 6.7b 41.03 0.13 $\pm$ 0.06 31.65 0.22 $\pm$ 0.12
DeepSeekCoder-Base 6.7b 32.62 1.01 $\pm$ 0.42 27.76 1.25 $\pm$ 0.98
CodeLlama-Instruct 7b 32.83 0.31 $\pm$ 0.15 23.49 0.36 $\pm$ 0.26
EditCoder 1.3b 26.67 0.14 $\pm$ 0.09 21.43 0.20 $\pm$ 0.12
DeepSeekCoder-Instruct 1.3b 26.22 0.32 $\pm$ 0.18 17.27 0.32 $\pm$ 0.13
DeepSeekCoder-Base 1.3b 17.90 0.69 $\pm$ 0.42 11.76 2.79 $\pm$ 2.29
Table 3: Evaluation results of close and open-access models on CanItEdit. We report the pass@1 and ExcessCode metrics for both the descriptive and lazy prompts as well as the size of the model if available.
In this section, we evaluate the performance of various open and closed models on the CanItEdit benchmark, as well our fine-tuned models.
@@
We run the open-access models using HuggingFace Transformers (Wolf et al., 2020) and vLLM (Kwon et al., 2023). We use the following hyperparameters for all inference experiments: $2048$ maximum new tokens, temperature $0.2$, and top-$p$ sampling cutoff of $0.95$. Following Cassano et al. (2023b), we sample 20 completions for each problem. We run all tests in a Docker container to mitigate the risk of malicious code execution.
@@
We evaluate several state-of-the-art models of varying sizes, fine-tuning some of them to build EditCoder. We group the models into two categories: open, models which we have access to their weights, and closed, models which we do not. We prompt each model with their recommended prompt template. The specific templates used appear in § A.6. The full list of models and their sizes appears in Table 3.
5.1 Evaluation Metrics
We employed two metrics to assess model performance: pass@k assesses functional correctness, and ExcessCode assesses conciseness and precision of code edits.
  • pass@k is the likelihood that at least one successful edit was made from $k$ attempts, as assessed by the test suite. In this section, we show results only for pass@1, and evaluation results for pass@10 and pass@100 with higher temperatures can be found in § A.2.
  • ExcessCode evaluates the presence of unnecessary code changes, as indicated by the fraction of changed lines not covered by the test suite. We calculate this metric by averaging the mean line coverage for passing completions across all problems, omitting those with no successful completions. The Python code used to calculate this metric is found at § A.1. We additionally report the standard error of the mean for this metric.
5.2 Results with Existing Models
@@
Closed source models outperform open source models. Our evaluation indicates a significant performance disparity between open and closed models. GPT-4, despite not being specifically trained on code-related tasks, surpasses DeepSeekCoder-Instruct 33b – the leading open source model – by an average of 13% in pass@1 for both descriptive and lazy tasks. DeepSeekCoder-Instruct stands out as the only open model exceeding any closed model’s performance, surpassing GPT-3.5-Turbo for descriptive prompts.
DeepSeekCoder-Instruct utilizes an undisclosed instruction-tuning dataset, therefore direct comparisons with other open models may not be entirely fair. In contrast, Mixtral-Instruct (Jiang et al., 2024), comparable to OpenAI models in its general instruction-following training focus, significantly lags in performance against both closed models and open models specialized in code generation tasks. Lastly, CodeLlama-Instruct-70b, ranked second in instruction-following capabilities, was developed using a dataset of examples generated by Llama 2 70b with a larger focus on code generation tasks. This may explain its superior performance compared to Mixtral-Instruct.
Descriptive prompts yield better performance than lazy prompts. Descriptive prompts result in a 8.68 absolute increase on average in pass@1 compared to lazy prompts, which may be the result of more detailed information and additional pointers in descriptive problems. Lazy instructions generally lead to lower ExcessCode for larger models. This may be indicative of lazy instructions introducing less noise in the task, as there are less tokens to attend to in the prompt given.
Larger models perform better than smaller models. Model size correlates positively with pass@1, and negatively with ExcessCode, indicating that larger models are more adept at precise edits to code. This pattern is most clearly seen in the evaluation results of DeepSeekCoder-Base, StarCoderBase, and StarCoder2, where a steady increase in performance is seen along with an increase in model size (§ A.5).
Models pre-trained on commits are better at code editing. Among open models, StarCoder is pre-trained on GitHub commits, while StarCoder2 focuses on GitHub issues. StarCoder outperforms similar-sized DeepSeek and CodeLlama models in our benchmark, despite their superior code generation capabilities. OctoCoder, a StarCoder-based model fine-tuned on instructions (Muennighoff et al., 2023), shows lower pass@1 performance, suggesting instruction fine-tuning on commit-based models may reduce code editing efficacy. Conversely, StarCoder2, exchanging commits with issues in its training, improves in editing tasks with larger models (§ A.5). This may be attributed to extended training and modern architecture rather than data source shift.
5.3 Results after Fine-Tuning on Commits
Training Dataset Metrics
Name #Tokens #Items pass@1 ExcessCode
Commits2023FT+EditPackFT 74M 46,274 43.81 0.34 $\pm$ 0.15
Commits2023FT 62M 24,129 41.88 0.33 $\pm$ 0.17
Commits2023Raw25k 62M 25,000 38.86 0.3 $\pm$ 0.12
EditPackFT 12M 22,602 41.6 0.26 $\pm$ 0.14
Table 4: Ablation results of training DeepSeekCoder-6.7b-Base on different datasets and evaluating on CanItEdit. We show the total number of tokens and items in each dataset, as well as the pass@1 and ExcessCode metrics for both the descriptive and lazy prompts aggregated across all problems. The reported sizes of the datasets are after deduplication.
@@ 5.3 Results after Fine-Tuning on Commits
In addition to evaluating existing open models, we also fine-tuned pre-trained DeepSeek models (§ 4) to build EditCoder, which we now evaluate.
@@
Optimal Dataset: Commits2023FT+EditPackFT. In finding the best training dataset for DeepSeekCoder-6.7b-Base, the base model for EditCoder, various ablation datasets were tested. Results in Table 4 show Commits2023FT+EditPackFT is the top performer for both descriptive and lazy instructions. The dataset’s larger size and diverse data types, including varied commits, edits, and instructions, likely contribute to its superior performance.
Fine-tuning on open commits can significantly improve code editing performance. EditCoder-33b surpasses all open models in pass@1 for both descriptive and lazy instructions types, showing an overall $10.7\%$ increase in pass@1 and a notable decrease in ExcessCode compared to its base model, DeepSeekCoder-Base-33b. Additionally, we see a substantial increase in pass@1 for every iteration of EditCoder over its corresponding base model, with the largest improvement being a $45.1\%$ increase at 6.7b.
Both EditCoder-33b and EditCoder-6.7b outperform GPT-3.5-Turbo in pass@1 for descriptive instructions, with EditCoder-33b also matching GPT-3.5-Turbo in for lazy ones. In higher temperature scenarios (§ A.2), EditCoder-33b beats GPT-3.5-Turbo in both instruction types for pass@10 and pass@100, and even surpasses GPT-4 in pass@100 for descriptive instructions. Analysis in § A.3 shows EditCoder excels in corrective changes but is less effective in perfective changes. Further analysis in § A.4 shows that EditCoder-33b outperforms all models, including GPT-4, in simple single-function bug fixes, and retains synthesis capabilities. This demonstrates the effectiveness of targeted fine-tuning on code editing datasets, addressing the distinct needs of instructional code editing compared to general code generation.
6 Conclusion